### v1

# Modification Takes Courage: Seamless Image Stitching via Reference-Driven Inpainting

Abstract
Current image stitching methods often produce noticeable seams in challenging scenarios such as uneven hue and large parallax. To tackle this problem, we propose the Reference-Driven Inpainting Stitcher (RDIStitcher), which reformulates the image fusion and rectangling as a reference-based inpainting model, incorporating a larger modification fusion area and stronger modification intensity than previous methods. Furthermore, we introduce a self-supervised model training method, which enables the implementation of RDIStitcher without requiring labeled data by fine-tuning a Text-to-Image (T2I) diffusion model. Recognizing difficulties in assessing the quality of stitched images, we present the Multimodal Large Language Models (MLLMs)-based metrics, offering a new perspective on evaluating stitched image quality. Compared to the state-of-the-art (SOTA) method, extensive experiments demonstrate that our method significantly enhances content coherence and seamless transitions in the stitched images. Especially in the zero-shot experiments, our method exhibits strong generalization capabilities. Code: https://github.com/yayoyo66/RDIStitcher
Figure 1: Different solutions to image fusion in image stitching. Our method reformulates the fusion and rectangling tasks as a reference-driven inpainting model. By boldly using a larger fusion modification area compared to UDIS++ [25] and SRStitcher [48], and applying a stronger modification intensity than UDIS [22] and UDIS++, we achieve a significant advancement in seamless image stitching, particularly in the challenging scenarios involving uneven hue and large parallax. Notice that UDIS and UDIS++ are three-stage architecture methods requiring additional rectangling models to complete the stitching process. Therefore, the rectangling areas for these methods are left blank in the figure.
1 Introduction
Image stitching is a fundamental problem in computer vision, which aims to seamlessly integrate multiple images captured from different perspectives into a wide field-of-view composite image [19, 42, 9]. Image fusion is the core stage of the image stitching pipeline and focuses on combining aligned images without visible seams or artifacts. However, as illustrated in Fig.1, this stage encounters two major challenges: (1) Uneven hue. Due to variations in atmospheric lighting conditions and camera settings, images taken from different viewpoints of the same scene may display inconsistent hues. When fusing images with significant hue differences, visible seams will likely appear in the stitched image. (2) Large parallax. Large parallax refers to the significant difference in the relative positions of objects in a scene when captured from different viewpoints. Existing homography-based registration methods [21, 20, 16] struggle to accurately align images in large parallax scenes, leading to noticeable artifacts and misalignment of content in the stitched images.
Current image fusion methods can be divided into three categories, including reconstruction-based (recon-based), seam-based, and inpainting-based. Recon-based methods [21, 24, 22] use pixel-by-pixel reconstruction to smooth the fused image, effectively handling scenes with uneven hues. However, in large parallax scenarios, recon-based methods can introduce notable artifacts, which degrade image quality. Additionally, seam-based methods [25, 7, 14] work by identifying optimal seams for image fusion. Nevertheless, these methods heavily rely on the assumption that perfect seams exist, which often fails to hold true in uneven hue and large parallax scenarios. Finally, the inpainting-based method [48] proposes to modify the fusion area to improve fusion effects. Unfortunately, the existing method is conservative in the size selection of modification fusion areas, so it is difficult to deal with uneven hue and large parallax scenarios. To detail the advantages and disadvantages of the three methods in different challenge scenarios, we conducted a small user experience survey presented in the Fig.2.
To address the limitations of current methods, we propose a key principle: Modification takes courage, including area size and intensity. We develop the RDIStitcher, which utilizes a larger modified area for fusion than previous seam-based and inpainting-based methods. Compared to recon-based methods, RDIStitcher applies stronger modification intensity.
However, larger and stronger modifications come at a cost, which is introducing more content instability into the stitched image. Due to the shortage of labeled data, SRStitcher is unable to train the model and can only maintain content consistency before and after stitching by limiting the size of the fusion modification area. Therefore, to implement RDIStitcher, we propose a self-supervised training method. Specifically, we leverage pre-knowledge from an unlabeled image stitching dataset [22] by using a pre-trained registration model to generate pseudo-stitching images based on single-view images. Subsequently, we apply a large-scale pre-trained T2I diffusion model [2] to learn the restoration of the single-view image based on the pseudo-stitching images, effectively teaching the model the new concept of stitching. Our experiments demonstrate that this self-supervised training method achieves remarkable results with high generalization ability, requiring only a few parameters to fine-tune the T2I model.
Figure 2: A user experience survey of the recon-based method UDIS [22], the seam-based method UDIS++ [25], and the inpainting-based method SRStitcher [48] on uneven hue and large parallax scenes. Please see the Supplementary Material for more details.
After designing the model and training method, the final task is to measure the stitched image quality in challenging scenarios without ground truth. The previous works UDIS [22] and UDIS++ [25] rely solely on small-scale user evaluations to assess stitched image quality, which is costly and lacks comprehensiveness. In addition, SRStitcher [48] introduces the No-Reference Image Quality Assessment (NR-IQA) metrics. However, existing NR-IQA techniques [30, 34] have significant flaws when applied to assessing stitched image quality, especially in detecting fine-grained stitching issues [48]. To address these challenges, we develop assessment methods based on MLLMs for stitched images, including the Single-Image Quality Score (SIQS) and the Multi-Image Comparative Quality Score (MICQS).
We summarize the main contributions as follows:
  • We reformulate the fusion and rectangling tasks as a reference-driven inpainting model. This model achieves remarkable stitching effects in challenging scenarios including uneven hues and significant parallax, while preserving the original structure and content of the input images. (Sec.3.1 and Sec.3.2)
  • We introduce a self-supervised training method that enables RDIStitcher to be trained without the need for labeled data. This method requires fine-tuning only a small number of parameters in a large-scale pre-trained T2I model, resulting in low hardware requirements. To our knowledge, this is the first unsupervised training method for the rectangling problem. (Sec.3.3)
  • We propose quality metrics for assessing stitched images by MLLMs. By incorporating the MLLMs into the image stitching domain for the first time, we offer a pioneering research perspective for the evaluation of stitched images. (Sec.4.2)
2 Related Work
2.1 Image Stitching
The previous learning-based image stitching methods typically follow a three-stage architecture: (1) Registration stage: In this stage, works [21, 20, 16] estimate the homography between the input images using networks. Then, they warp the input images to a unified coordinate system to obtain aligned images. (2) Fusion stage: This stage focuses on fusing the aligned images into a seamless composite image. Existing methods can be divided into recon-based methods [21, 24, 22] and seam-based methods [25, 7, 14]. (3) Rectangling stage: The objective of this stage is to transform irregularly edged stitched images into regular rectangular formats. Current rectangling methods [23, 36, 38, 46] are all supervised and rely on labeled datasets.
Beyond the three-stage architecture, the inpainting-based method SRStitcher [48] simplifies the image stitching pipeline by integrating fusion and rectangling into a unified model. SRStitcher is implemented based on a large-scale pre-trained generative model, requiring no additional training or fine-tuning. However, this methodology implies that the performance of SRStitcher is entirely dependent on the capabilities of the pre-trained model. Consequently, SRStitcher adopts a highly conservative strategy in managing the fusion region size to avoid excessive alterations, but this results in noticeable seams in challenging scenarios. Furthermore, the generalization capabilities of SRStitcher are inherently limited by the pre-trained model, making it difficult to extend to domains such as medical imaging [5, 1] and remote sensing [35, 33].
Our method builds on the framework established by SRStitcher and further advances it by proposing a reference-driven inpainting model for fusion and rectangling. Additionally, we implement a self-supervised training method that is specifically designed for this new modeling paradigm. This enhancement increases the scalability and adaptability of the inpainting-based image stitching pipeline. Furthermore, our method represents a significant advancement in addressing the rectangling problem, as it offers the first training method that does not require labeled data.
2.2 Reference-Driven Inpainting
Image inpainting aims to fill missing image regions with coherent results [8, 41, 17]. Reference-driven inpainting [39, 47, 31, 6] is a new subfield of image inpainting research that focuses on filling missing regions leveraging content from reference images. Although some reference-driven inpainting methods [47, 31] show potential for applications in image stitching, they have not developed a complete training pipeline or conducted comprehensive experimental validation.
In contrast, our method introduces the first image stitching pipeline based on reference-driven inpainting, encompassing model design, training, and validation, as well as introducing the novel evaluation metrics. Our method employs the input and output construction strategy of LeftRefill [6], but differs significantly in terms of the training method, model structure, and the number of parameters.
3 Method
3.1 Problem Definition
Figure 3: The framework of RDIStitcher. (a) Training. For the sake of clarity in the presentation, the input images and masks are simplified. (b) Inference. Details on the specific input structure. (c) Data processing in self-supervised. Details on the self-supervised training method.
Given a pair of input images: a reference image and a target image, defined as $I_{r},I_{t}\in\mathbb{R}^{H\times W}$, where $H$ and $W$ are the height and width, respectively. Similar to SRStitcher [48], the image stitching pipeline designed in this paper is a two-stage architecture.
In the first stage, the homography matrix $\mathcal{H}$ between the input images is estimated by a registration model. Then, the $I_{r}$ is aligned with the $I_{t}$ by a warp function $\mathcal{W}(\cdot)$ based on $\mathcal{H}$. The process of obtaining the aligned images $I_{wr}$ and $I_{wt}$ can be expressed as Eq.1.
$I_{wr},I_{wt}=\mathcal{W}(I_{r},\mathcal{H}),\mathcal{W}(I_{r},\mathbb{I}),$ (1)
where, $\mathbb{I}$ is the identity matrix. The masks $M_{wr}$ and $M_{wt}$ corresponding to the aligned images can also be obtained by the Eq.1, except that the input images is replaced by the all-ones matrixes of the same size.
In the second stage, we reformulate the fusion and rectangling problems as a reference-driven inpainting model. Given a target image $I_{wt}$ and its corresponding mask $M_{wt}$, the objective is to train a model $\epsilon_{\theta}(\cdot)$ that inpaints the image $I_{wt}$ in accordance with the mask $\overline{M}_{wt}=1-M_{wt}$ to obtain the stitched image $I_{s}$, where the inpainting process is constrained by the content of the reference image $I_{wr}$, as represented in Eq.2.
$I_{s}=I_{wt}\odot M_{wt}+\epsilon_{\theta}(I_{wr},I_{wt})\odot\overline{M}_{wt},$ (2)
where, $\odot$ denotes pixel-wise multiplication. Different from SRStitcher [48], RDIStitcher employs the $\mathcal{W}(\cdot)$ from UDIS [22] instead of UDIS++ [25]. Because the $\mathcal{W}(\cdot)$ of UDIS++ occasionally introduces huge local distortions into the aligned images, which could potentially disrupt the fitting process that relies on undistorted images. In addition, we select the image warped by $\mathbb{I}$ as the target image, it is also a consideration of the degree of image distortion.
3.2 Framework of RDIStitcher
We propose the introduction of a large-scale pre-trained T2I model to enhance the generalization capabilities in the design of $\epsilon_{\theta}(\cdot)$. Furthermore, we develop a training method based on self-supervised paradigms to improve the practical feasibility. In general, the framework of RDIStitcher is illustrated in Fig.3.
[…]
To enable the T2I model to perform specific image stitching tasks, we incorporate the unique identifier from DreamBooth [29] into our method. Specifically, we define a special text prompt, denoted as $\mathcal{P}$, which guides the model in carrying out reference-driven inpainting operations. During training, we use the Low-Rank Adaptation (LoRA) [13] to fine-tune both the text encoder and the U-Net of the T2I model, effectively reducing memory usage.
3.3 Self-Supervised Training Method
A significant challenge in training $\epsilon_{\theta}(\cdot)$ is the extreme scarcity of available data, particularly labeled datasets for image stitching [22]. The lack of labeled data makes it difficult to directly apply mainstream supervised learning paradigms. To overcome this challenge, we turn to the self-supervised learning paradigm, which aims to generate meaningful training signals by exploiting the inherent structure or relationships present in the input data, without the need for explicit annotations [11, 26].
To learn image stitching concepts, we hypothesize: We can use a single-view image to simulate the pseudo-stitched images and then train a model to restore the original image based on the pseudo-stitched images. Through this process, the model can establish an understanding of the underlying stitching relationships in the data. Subsequently, we can leverage the strong generalization capabilities of the large-scale T2I model to generalize this knowledge to real stitching data.
As shown in Fig.3 (c), we collect the stitching mask distribution from a large-scale image stitching dataset UDIS-D [22] by a pre-trained registration model [22]. The pre-knowledge mask distribution is defined as $\mathsf{M}_{N}=\{(M^{1}_{wr},M^{1}_{wt}),(M^{2}_{wr},M^{2}_{wt}),...,(M^{N}_{wr},M^{N}_{wt})\}$, where $N$ is the number of samples in UDIS-D training set $D_{train}$. Then, suppose $I_{sg}$ is a single-view image in the $D_{train}$ and $(M^{i}_{wr},M^{i}_{wt}),i\in[1,N]$ is a random mask set from $\mathsf{M}_{N}$, the pseudo-stitching reference and target images $\tilde{I}_{wr},\tilde{I}_{wt}$ can be obtained by Eq.3.
$\tilde{I}_{wr},\tilde{I}_{wt}=I_{sg}\odot M^{i}_{wr},I_{sg}\odot M^{i}_{wt}.$ (3)
However, the pseudo-stitching images produced by the above method are inherently perfectly correct and lack challenging scenarios. To improve the robustness, we design two data augmentation strategies to simulate color differences of uneven hue conditions and misalignments of large parallax conditions.
Data augmentation for uneven hue. To simulate color differences, we add random color jitter $\mathsf{ColorJitter}(\cdot)$ [28] with probability $p_{cj}$ to the pseudo-stitching reference image $\tilde{I}_{wr}$, as formulated in Eq.6.
$\tilde{I}_{wr}=\left\{\begin{array}[]{l}\mathsf{ColorJitter}(\tilde{I}_{wr},\mathbf{E}_{cj})\quad\text{with }p_{cj},\\ \tilde{I}_{wr}\qquad\qquad\qquad\quad\text{with }1-p_{cj},\\ \end{array}\right.$ (6)
where, $\mathbf{E}_{cj}=\{e_{b},e_{c},e_{s},e_{h}\}$ are the hyper-parameters of the brightness, contrast, saturation and hue adjustments.
Data augmentation for large parallax. The registration errors induced by large parallax are manifested differently across various warping methods. In UDIS++ [25], these errors result in pronounced local distortions, while in UDIS [22], they lead to content misalignment near the seams. Obviously, content misalignment is comparatively simpler to simulate, so we select the $\mathcal{W}(\cdot)$ of UDIS as the foundation for designing RDIStitcher. To simulate misalignments, a random affine transformation $\mathsf{Aff}(\cdot)$ with probability $p_{at}$ is applied to the $\tilde{I}_{wr}$ and its corresponding mask $M^{i}_{wr}$, as demonstrated in Eq.9.
$\tilde{I}_{wr}=\left\{\begin{array}[]{l}\mathsf{Aff}(I_{sg},\mathcal{M}(t_{x},t_{y}))\odot M^{i}_{wr}\quad\text{with }p_{at},\\ I_{sg}\odot M^{i}_{wr}\qquad\qquad\qquad\quad\text{with }1-p_{at},\\ \end{array}\right.$ (9)
where, $\mathcal{M}(\cdot)$ is a $2\times 3$ affine transformation matrix whose horizontal $t_{x}\in[-(W^{*}-x_{max}),x_{min}]$ and vertical $t_{y}\in[-(H^{*}-y_{max}),y_{min}]$ translations are random values, $W^{*}$ and $H^{*}$ are the width and height of $I_{sg}$. Also, $x_{max}$, $x_{min}$, $y_{max}$ and $y_{min}$ are coordinates of the smallest enclosing rectangle of $M^{i}_{wr}$.
Coarse rectangling and gradient mask. The coarse rectangling and gradient mask are originally introduced by SRStitcher [48]. Coarse rectangling aims to reduce the likelihood of generating abnormal content in the rectangling regions by incorporating weak priors. And, the gradient mask is utilized to smooth the seam regions. However, SRStitcher claims that coarse rectangling introduces a local blurring side effect that degrades image quality. To address this issue, we propose a self-supervised training method that allows the model to establish the relationship between the weak priors and the concrete images, effectively mitigating the local blur problem. Additionally, we reconstruct the gradient mask based on the RDIStitcher input mask to enhance image cohesion in the seam regions. Specifically, Eq.10 details the process of obtaining the coarse rectangling image $\tilde{I}^{cr}_{wr}$ by the alexandru telea algorithm $\mathsf{Telea}(\cdot)$ [32].
$\tilde{I}^{cr}_{wr}=\mathsf{Telea}(\tilde{I}_{CF},M^{i}_{wr}\vee M^{i}_{wt},R),$ (10)
where, $\tilde{I}_{CF}=\tilde{I}_{wr}+\tilde{I}_{wt}\odot(1-(M^{i}_{wr}\&M^{i}_{wt}))$, $\vee$ and &\&& denote the bitwise OR and AND operators, and $R$ is the hyper-parameter to control the radius. Furthermore, We use Eq.11 to generate the seam gradient target mask $\mathfrak{M}^{i}_{wt}$.
$\mathfrak{M}^{i}_{wt}=\mathsf{Blur}(\mathsf{Dilation}(M^{i}_{wt},K_{d}),K_{g}),$ (11)
where, $\mathsf{Dilation}(\cdot)$ is the dilation operation [27] with kernel $K_{d}$, and $\mathsf{Blur}(\cdot)$ is the Gaussian blur operation [27] with kernel $K_{g}$.
Objective function. RDIStitcher is based on Denoising Diffusion Probabilistic Model (DDPM) [12], where the $\epsilon_{\theta}(\cdot)$ is transformed into a noisy prediction model. Suppose the latents $\mathbf{x}_{0},\mathbf{x}_{1},...,\mathbf{x}_{T}$ are derived by adding Gaussian noise to the original data $\mathbf{x}_{0}\sim q(\mathbf{x}_{0})$ with $T$ steps, where $\mathbf{x}_{0}=\mathsf{ImageEncoder}(\mathsf{Concat}(\tilde{I}^{cr}_{wr},I_{sg}))$. The objective function $\mathcal{L}$ of $\epsilon_{\theta}(\cdot)$ in RDIStitcher is defined as in Eq.12.
$\mathcal{L}=\mathbb{E}_{\vartheta}\left[\left\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},t,\mathcal{P},\mathbf{M}^{i},\overline{\mathbf{M}^{i}}\odot\mathbf{x}_{0})\right\|^{2}\right],$ (12)
where, $\mathbf{M}^{i}=\mathsf{Concat}(M_{0},\mathfrak{M}^{i}_{wt})$, $\overline{\mathbf{M}^{i}}=1-\mathbf{M}^{i}$, $t\in[1,T]$, $\vartheta=\{\mathbf{x}_{0},t,\epsilon,\mathbf{M}^{i}\}$, and $\epsilon\sim\mathcal{N}(0,1)$.
4 Experiment
4.1 Experimental Setup
Dataset. Most of the experiments in this section are conducted on the unsupervised dataset UDIS-D [22]. To the best of our knowledge, UDIS-D is the only publicly available dataset in learning-based image stitching. To provide a more comprehensive evaluation, we conduct cross-dataset experiments in a zero-shot manner on three traditional datasets, including the APAPdataset [40], the REWdataset [15], and the SPWdataset [18]. However, since these traditional datasets are small and contain only tens of samples, performing quantitative evaluations on them would not produce meaningful results. Therefore, this work focuses on a qualitative evaluation to demonstrate the differences in the generalization capabilities on traditional datasets.
Baseline method. Selecting appropriate baseline methods is a little difficult, because our method is the first to propose a fusion and rectangling unified model that can be trained on the unsupervised image stitching dataset, a groundbreaking advancement in the image stitching. As mentioned in Sec.2.1, UDIS [22] and UDIS++ [25] can only perform up to the fusion stage and require integration with another rectangling model to complete the entire image stitching pipeline. However, the SOTA rectangling models [23, 36, 38, 46] are all supervised and can only be trained on a task-specific dataset DIR-D [23], rather than the real-world image stitching dataset UDIS-D. This limitation makes it difficult to compare these methods with RDIStitcher under relatively fair conditions. To address this issue, we propose a variant of RDIStitcher, called RDIStitcher-R, which works only as a rectangling model trained with the same settings as RDIStitcher. With this, we can combine the RDIStitcher-R with UDIS and UDIS++ to establish baselines. In summary, the baselines include the SOTA recon-based method UDIS with RDIStitcher-R (UDIS+R) and the SOTA seam-based method UDIS++ with RDIStitcher-R (UDISplus+R). The implementation details of RDIStitcher-R can be found in the Supplementary Material.
In addition, the baselines also include the SOTA inpainting-based method SRStitcher [48]. Although the SRStitcher method does not involve any training or fine-tuning, which makes the comparison somewhat less fair, we believe it is necessary to include the results of this most important related work. The experimental results demonstrate that our method significantly outperforms SRStitcher. This result represents a noteworthy advancement in the inpainting-based image stitching framework.
Finally, we also implement our self-supervised training method on LeftRefill [6] as a baseline.
Implement detail. RDIStitcher is fine-tuned on a pre-trained T2I model [2] with a batch size of 4, input image size 1024$\times$512, LoRA rank 8, learning rate 2e-4 for the U-Net, learning rate 4e-5 for the text encoder, and 10K iterations on two NVIDIA 4090 GPUs. And, all results of RDIStitcher described in the experiment are based on the pre-aligned images made by UDIS [22]. For hyper-parameters, the $\mathbf{E}_{cj}$ and $p_{cj}$ of Eq.6 are set to $\{0.2,0.2,0.2,0.1\}$ and 0.25; The $p_{at}$ of Eq.9 is set to 0.25; The $R$ of Eq.10 is set to 3; The $K_{d}$ and $K_{g}$ of Eq.11 are set to 10 and 15. For inference, the inference step is set to 50. More details can be found in Supplementary Material.
4.2 MLLMs-based Metric
Table 1: An example of SIQS evaluated by Qwen-VL-Max.
Evaluation Prompt
I need you to become a stitched image quality assessment evaluator. The evaluation process should be as objective and impartial as possible, giving specific ratings and reasons, including seam, brightness transition, distortion, clear and abnormal content, each aspect 2 points.
[The full example is in the Supplementary Material.]
Reason
- Seam The image shows slightly visible boundaries, particularly noticeable in the sky and along the edges of the buildings. (1 points).
- Brightness transition The brightness transition in the image is relatively smooth. (2 points).
- Distortion There are no noticeable distortions in the image. The straight lines of the buildings and the street appear to be accurate. (2 points).
- Clear The buildings, palm trees, and streets are all distinguishable, and the details are visible. (2 points).
- Abnormal content The image contains some unnatural elements, such as the seams. (1 points).
Score
The overall impression is that the image is a stitched panorama with noticeable flaws. (8 points).
Assessing the quality of stitched images on unlabeled datasets has long been a challenging problem in the field. Previous works UDIS [22] and UDIS++ [25] rely solely on small-scale user evaluations to measure stitched image quality, which lacks comprehensiveness. Additionally, SRStitcher [48] proves that the mainstream NR-IQA metrics [30, 34] struggle to accurately evaluate stitched image quality, particularly in evaluating fine-grained stitching issues such as uneven hue and large parallax.
Quality measurement challenges are not unique to image stitching. After surveying the broader research landscape, we observe that many emerging domains, such as personalization generation [3, 43], face similar measurement difficulties. In response, some studies [37, 44, 45] have proposed leveraging MLLMs to evaluate the quality of unlabeled images, leading to significant advancements in the field.
Motivated by these developments, we ask: Can MLLMs be leveraged as effective evaluators to measure the quality of stitched images? To explore this, we propose the MLLMs-based stitched image evaluation metrics, including the Single-Image Quality Score (SIQS) and the Multi-Image Comparative Quality Score (MICQS).
To clarify the metrics, an example of SIQS is presented in Table.1, conducted by Qwen-VL-Max [4], denoted as SIQS-Q. In addition, Table.2 provides the MICQS evaluation comparing two stitched images, also performed by Qwen-VL-Max, denoted as MICQS-Q. To mitigate potential biases associated with relying on a single model, we introduce the model GLM-4V-Plus [10] as an additional evaluator (SIQS-G, MICQS-G).  [^1] We do not employ widely used MLLMs such as GPT-4V or Claude-3-Haiku, due to the limited access..
Table 2: An example of MICQS evaluated by Qwen-VL-Max.
[Uncaptioned image]          [Uncaptioned image]
Evaluation Prompt
I need you to become a stitched image quality assessment evaluator. Compare the two input stitched images and choose which one you think is better. There are two choices: image 1 (left) is better, or image 2 (right) is better.
[The full example is in the Supplementary Material.]
Reason
- Seam Image 2 has a smoother transition between the stitched sections compared to image 1, where the seam is more noticeable.
- Brightness transition Image 2 has a more consistent brightness level across the entire image, while image 1 shows a noticeable difference in brightness.
Conclusion
Image 2 (right) is better.
Other metrics. In addition to the MLLMs-based image quality metrics, we also introduce the Content Consistency Score (CCS) from SRStitcher [48] and the real user evaluation from UDIS [22] with the same settings.
4.3 Quantitative Evaluation
Single-image evaluation. Table.3 presents the comparative experimental results of SIQS-Q, SIQS-G, and CCS on the UDIS-D test set with 1,106 sample pairs. Our method surpasses other methods in terms of both quality and content consistency.
Multi-image comparative evaluation. This method evaluates the stitched image quality in a contrastive manner, including MLLMs-based evaluators and real user evaluators. The results are presented in the Fig.4, which illustrates that our method achieves significantly higher votes, clearly outperforming the methods.
Table 3: Single-image evaluation results. We use five seeds and then take the mean and standard deviation. The best performer (prioritize the mean) is highlighted by red.
Method SIQS-Q SIQS-G CCS (%)
UDIS+R 9.40$\pm$ 0.08 9.27$\pm$ 0.05 89.64$\pm$ 0.96
UDISplus+R 9.49$\pm$ 0.05 9.33$\pm$ 0.04 90.17$\pm$ 0.88
SRStitcher 9.28$\pm$ 0.07 8.44$\pm$ 0.09 91.15$\pm$ 0.52
LeftRefill 9.23$\pm$ 0.07 8.51$\pm$ 0.13 83.98$\pm$ 1.85
RDIStitcher (Ours) 9.54$\pm$ 0.07 9.39$\pm$ 0.05 91.23$\pm$ 0.79
Figure 4: Multi-image comparative evaluation results. We decide not to give MLLMs-based evaluators the both option as we discover that they consistently favor both good.
4.4 Qualitative Evaluation
The qualitative evaluation results are shown in Fig.5, highlighting the significant advantages of our method across various challenging scenarios. Notably, in the zero-shot experiments, our method demonstrates more obvious advantages, which verifies the generalization capability hypothesis proposed in Sec.3.3.
Figure 5: Qualitative evaluation results. The upper half of the dotted line displays the results on UDIS-D, and the lower half is the results on traditional datasets. We highlight the areas with significant seams and errors using the local magnification and arrows. Special emphasis is placed on the error regions for the LeftRefill method, providing a thorough analysis of its performance limitations. Notice that the last example represents a hybrid challenge with uneven hue, large parallax, and zero-shot conditions. In this highly complex scene, our method shows exceptional performance, markedly surpassing that of previous methods. More results can be found in the supplementary Material.
4.5 Discussion
Limitation of LeftRefill. The results presented in Table.3 and Fig.5 indicate that LeftRefill [6] performs significantly worse than other methods. By observing and analyzing the experimental results, we believe that LeftRefill struggles to accurately differentiate between fusion and rectangling tasks. As a result, it inpaints the coarse rectangling image into the stitched image, leading to significant errors. While LeftRefill achieves reference-driven inpainting with minimal parameter adjustments (hundreds of kilobytes), image stitching is a complex mixed task that includes both fusion and rectangling. This complexity poses challenges for LeftRefill. Therefore, in developing RDIStitcher, we use more fine-tuned parameters (dozens of megabytes) and a different fine-tuned network structure than LeftRefill. Although the saved LoRA weights are slightly larger, we believe that this difference does not significantly affect the usability of our method. In addition, despite the increased number of adjusted parameters, the memory required for training our method is much lower than that for LeftRefill, which may be caused by some operations required by LeftRefill on the attention modules.
Ablation study. Our method is designed to be inherently interpretable at each module, which makes ablation experiments less critical. Therefore, ablation experiments are not included in the main paper. The Supplementary Material offers additional information that may interest readers, including different hyper-parameter configurations and the robustness of our method with different random seeds.
5 Conclusion and Future Work
In this paper, we propose an image stitching model based on reference-driven inpainting, which achieves remarkable seamless stitching results on challenging scenes. Extensive experiments demonstrate the effectiveness of our model, particularly highlighting its strong generalization performance in zero-shot scenarios. However, through a comprehensive analysis of the experimental results, we believe that there are three critical open issues to be addressed in the future:
(1) Stability of content generated in rectangling regions. Table.3 shows that the standard deviation of content consistency in our method is relatively higher compared to that of SRStitcher. Because our method tries to provide specific images instead of local blur, which results in reduced stability. In the future, it may be possible to consider using a different loss to optimize the stability of rectangling regions.
(2) Accuracy of MLLMs-based stitched image quality metrics. Although the MLLMs-based metrics provide new insights into stitched image metrics, they still have limitations (see the Supplementary Material for details). However, we believe that MLLMs-based measurement is a promising research direction, and we are considering more further in-depth studies to introduce more comprehensive MLLMs and to develop prompts for comparison.
(3) Reliability of the generation-based methods. Some researchers express concern about the reliability of applying generation-based models to traditional vision problems, likening them to a lottery machine. However, we argue that, compared to previous methods, our method at least confirms that a big prize exists in the solution space. The future challenge is finding a more precise way to locate this big prize.
Supplementary Material
Appendix A Overview
In the supplementary material, we provide the following contents:
  • Source code for our work, including train, inference, and MLLMs-based metrics. Please see the code document;
  • Survey details of Fig.2 (Sec.B);
  • More Results of Qualitative Evaluations (Sec.D);
  • More Information of Ablation Studies (Sec.E);
  • More Details of MLLMs-based Metrics (Sec.F).
Appendix B Survey Details
We first select 50 challenging samples exhibiting uneven hue, large parallax, and mixed scenarios (containing both challenges) from the UDIS-D [22], APAPdataset [40], REWdataset [15], and SPWdataset [18]. After applying the UDIS [22], UDIS++ [25], and SRStitcher [48] to fuse these samples, we task four volunteer participants with rating the visual quality of the stitched images. Specifically, Volunteers provided scores of either good or bad based on two criteria: brightness and hue smoothness for uneven hues and content continuity and artifact severity in seam regions for large parallax. According to the evaluation results, methods achieving over 75% good ratings for a challenge receive a red smiley face, 50%-75% receive a yellow smiley face, 25%-50% a green face and less than 25% a blue sad face.
Appendix C More details of RDIStitcher-R
C.1 Implementation of RDIStitcher-R
RDIStitcher-R is a variant of RDIStitcher that only performs rectangling tasks, so its implementation is much simpler.
[…]
$\tilde{I}_{pf}=I_{sg}\odot(M^{i}_{wr}\vee M^{i}_{wt}).$ (13)
The coarse rectangling fusion image can be obtained by Eq.14.
$\tilde{I}^{cr}_{pf}=\mathsf{Telea}(\tilde{I}_{pf},M^{i}_{wr}\vee M^{i}_{wt},R).$ (14)
We also produce the gradient mask by Eq.15.
$\mathfrak{M}^{i}_{gm}=\mathsf{Blur}(\mathsf{Dilation}(M^{i}_{wr}\vee M^{i}_{wt},K_{d}),K_{g}).$ (15)
An overview of the architecture of RDIStitcher-R is shown in Fig.6.
Figure 6: The framework of RDIStitcher-R.
C.2 Ablation Study of RDIStitcher-R
RDIStitcher-R operates by fine-tuning the pre-trained T2I model to establish the relationship between the coarse rectangling image and the detailed image.
The Table.4 shows the ablations of USID+RDIStitcher-R (UDIS+R) based on the content consistency score CCS [48] on the UDIS-D test set. The comparisons include UDIS+stable diffusion inpainting model [2] without fine-tuning (UDIS+SD2) and UDIS+RDIStitcher-R without coarse rectangling (UDIS+RWCR). The results demonstrate that RDIStitcher-R effectively reduces the likelihood of abnormal content.
Table 4: Ablation study of RDIStitcher-R. The best performer is highlighted by red.
Method UDIS+SD2 UDIS+RWCR UDIS+R
CCS (%) 85.97$\pm$ 1.33 87.68$\pm$ 1.17 89.64$\pm$ 0.96
Appendix D More Results of Qualitative Evaluations
Due to space limitations, we only present a limited number of qualitative evaluation results in the main paper. Here, to further examine the generalization capabilities of the RDIStitcher, We provide additional qualitative results obtained on the APAPdataset [40], REWdataset [15], and SPWdataset [18].
The additional qualitative results on the APAPdataset [40] are presented in Fig.8. Note that in the fifth case, the stitched image produced by our method appears to exhibit some errors in the rendering of Chinese characters. We speculate that this is likely due to the fact that the underlying T2I model employed is not optimized for Chinese language processing. As a result, the model may not be well-equipped to handle the accurate representation of Chinese characters. This paper utilizes the Stable Diffusion 2 inpainting version, acknowledging that Stable Diffusion 3 has significantly improved character generation. However, since no open-source inpainting version is currently available for Stable Diffusion 3, we are unable to test our method on it. Based on research trends, we believe that the character generation issue will not remain a major bottleneck and will be effectively resolved soon.
Moreover, the additional qualitative results on the REWdataset [15] are shown in Fig.9. And the additional qualitative results on the SPWdataset [18] are shown in Fig.10. The scenarios of SPWdataset are a relatively simple, and it is evident that most methods demonstrate good performance.
Appendix E More Information of Ablation Studies
E.1 Different Hyper-parameter Configurations
LoRA Configurations LoRA [13] is considered one of the most influential technologies in recent years, with numerous studies exploring its mechanisms. Due to the limited space of the main paper, all parameters of LoRA are not introduced in detail. Here, we provide additional parameters: rank 8, alpha constant 16, dropout rate 0.1, and bias none. For more details, please refer to the source code we have provided.
We conducted an ablation test by varying the LoRA rank while keeping other parameters fixed. The results are shown in the Table.5. However, we quickly realized that this isolated exploration does not definitively demonstrate that a LoRA rank 8 represents the optimal parameter setting for our specific task. A comprehensive exploration of LoRA cnfigurations is complex and requires significant hardware resources, which is beyond the scope of this study. Therefore, we provide a reference set of LORA parameter values in the main paper, but have not pursued any further ablation analysis regarding the LORA settings.
Table 5: Ablation study of LoRA rank. The best performer is highlighted by red.
Method Rank 4 Rank 8 Rank 16
CCS (%) 90.54$\pm$ 0.92 91.23$\pm$ 0.79 89.35$\pm$ 0.83
Seed robustness We present the qualitative evaluation results with different random seeds in Fig.7. In cases of small parallax, the stitched image produced by our method maintains high consistency. For large parallax, differences in the fusion region are insignificant, and slight changes occur in the rectangling region. Importantly, our method does not introduce significant abnormal content, unlike the original T2I model. In zero-shot tests, the stitched images of uneven hue and small parallax scenes also maintain good consistency, with only minor changes in the rectangling region. In high challenge scenarios, while content changes in the stitched images are more pronounced, our method shows clear advantages in brightness smoothing and content connection compared to previous methods (see Fig.5 of main paper for comparison).
It is worth noting that the inherent characteristics of the generation-based model can lead to occasional instability, resulting in some failure cases. However, our experiments indicate that the failure rate of our method is very low and within an acceptable range. Future work will focus on further enhancing reliability .
Figure 7: Qualitative results on the different seeds.
Appendix F More Details of MLLMs-based Metrics
F.1 Full Example of MICQS Evaluation
Due to space limitations, the evaluation prompt and reasons of MICQS evaluation in the main paper is truncated. Here, Table.6 gives a complete example of MICQS evaluation based on Qwen-VL-Max [4].
Table 6: A full example of MICQS evaluated by Qwen-VL-Max.
[Uncaptioned image]        [Uncaptioned image]
Evaluation Prompt
I need you to become a stitched image quality assessment evaluator. Compare the input two stitched images, includes seam, brightness transition, distortion, clear and abnormal content. Choose which one you think is better. There are two choices: image 1 (left) is better, or image 2 (right) is better.:
Reason
- Seam Image 2 has a smoother transition between the stitched sections compared to image 1, where the seam is more noticeable.
- Brightness transition Image 2 has a more consistent brightness level across the entire image, while image 1 shows a noticeable difference in brightness between the sections.
- Distortion There are no noticeable distortions in the both images.
- Clear Image 2 appears clearer overall.
- Abnormal content Image 2 has less noticeable artifacts or abnormalities in the content.
Conclusion
Image 2 (right) is better.
F.2 Full Example of SIQS Evaluation
Due to space limitations, the evaluation prompt of SIQS evaluation in the main paper is truncated. Here, Table.7 gives a complete example of SIQS evaluation based on Qwen-VL-Max [4].
F.3 Experiment Settings of MLLMs-based Metrics
To evaluate the stitched images quality, we use the API provided by the MLLMs service. Occasionally, errors occurred for unknown reasons led to the absence of correct result returns. Therefore, for each test stitched image, we submit multiple test requests until a valid return value is obtained for evaluation.
F.4 Limitations of MLLMs-based Metrics
Prompt sensitivity Different MLLMs exhibit varying sensitivities to prompts. As depicted in Table.3, GLM-4V-Plus [10] scores significantly lower than other methods for SRStitcher [48] and LeftRefill [6], as blurred is a key scoring criterion. Although SRStitcher and LeftRefill may stitch images with numerous local blur instances, Qwen-VL-Max [4] is relatively insensitive, not displaying a significant score reduction. Therefore, the construction of prompts can significantly impact the evaluation performance of the MLLMs.
Score mismatch In a well-known case testing LLMs with 9.11 or 9.3 is bigger, most LLMs failed. In our experiment, we also observe a score mismatch phenomenon in MLLMs, where the sum of scores in each aspect can not equal the total score.
Concept limitation In the early design, we intended to include artifacts in the evaluation criteria of MLLMs-based metrics. However, incorporating Artifact-related prompt texts led to numerous false evaluations. Through the analysis of the results, we found that the MLLMs used in our paper actually does not know the concept of Artifact, resulting in significant misjudgments. This substantial error prompted us to replace the Artifact-related prompt texts with Abnormal Content. Although test results indicate that MLLMs struggle to assess what constitutes abnormal content, they at least avoid major misjudgments. It is evident that the current knowledge domain of MLLMs we utilize in the paper has considerable limitations in evaluating the quality of stitched images.
Unstable result Our empirical analysis reveals that when the same image is evaluated multiple times by MLLMs, the scores exhibit greater instability, especially GLM-4V-Plus [10]. In some cases, the scores can fluctuate with a variance of up to 4 points.
It is important to acknowledge that MLLMs represent a relatively nascent technology, with the existing public models having been available for testing only in the past two years. This limited timeframe has inevitably led to apparent defects and shortcomings in the current MLLMs. However, it is equally clear that both the research focus and the prevailing research trends strongly indicate the immense potential of MLLMs. Therefore, we believe that the use of MLLMs for stitched image quality evaluation represents a highly promising research direction, and its feasibility should not be hastily dismissed due to the defects present in the current MLLMs.
Figure 8: Additional qualitative results on the APAPdataset[40].
Figure 9: Additional qualitative results on the REWdataset [15].
Figure 10: Additional qualitative results on the SPWdataset [18].
Table 7: A full example of SIQS evaluated by Qwen-VL-Max.
Evaluation Prompt
I need you to become a stitched image quality assessment evaluator. The evaluation process should be as objective and impartial as possible, giving specific ratings and reasons, including seam, brightness transition, distortion, clear and abnormal content, each aspect 2 points.
1. Whether there are seams in the image (2 points).
score 2: the image is smooth without obvious boundaries or misalignment;
score 1: there are slightly visible boundaries in the image, but overall look well;
score 0: there are obvious borders or dislocations in the image, affecting the overall look and feel.
2. Whether there are brightness transitions in the image (2 points).
score 2: the brightness transition of image is smooth;
score 1: the light and shade changes in the image are a bit unnatural;
score 0: the light and shade changes in the image are very abrupt.
3. Whether there are distortions in the image (2 points).
score 2: no distortion in the image;
score 1: there are a few structural anomalies of straight lines in the image;
score 0: there are noticeably distortions, such as distorted pillar, brick, and building construction.
4. Whether the image is clear and blurred (2 points).
score 2: the image is clear, the details are visible, and there is no blur;
score 1: the resolution of the image is good, but slightly blurred;
score 0: the image is blurred and the details are not clear.
5. Whether the image is natural (2 points).
score 2: the image is natural with out abnormal content;
score 1: there are some places in the image that is not in harmony with the main content;
score 0: There are a lot of abnormal content in the image such as strange texture and non-semantic image.
Reason
- Seam The image shows slightly visible boundaries, particularly noticeable in the sky and along the edges of the buildings. (1 points).
- Brightness transition The brightness transition in the image is relatively smooth. (2 points).
- Distortion There are no noticeable distortions in the image. The straight lines of the buildings and the street appear to be accurate. (2 points).
- Clear The buildings, palm trees, and streets are all distinguishable, and the details are visible. (2 points).
- Abnormal Content The image contains some unnatural elements, such as the seams. (1 points).
Score
The overall impression is that the image is a stitched panorama with noticeable flaws. (8 points).